OpenStudy (a1234):

Find the exact value of the trig expression given that sin u = -5/13 and cos v = -4/5. u and v are in the second quadrant. cos(v - u) Is this -33/65?

OpenStudy (a1234):

@SolomonZelman

jimthompson5910 (jim_thompson5910):

how is angle u in the second quadrant but they state that `sin u = -5/13` ??

jimthompson5910 (jim_thompson5910):

sine is positive in Q2

OpenStudy (a1234):

Typo...only cos is negative. The sin is positive.

jimthompson5910 (jim_thompson5910):

so it says sin(u) = 5/13 ?

OpenStudy (a1234):

yes

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

-33/65 is incorrect

OpenStudy (displayerror):

You can use a difference of two angles formula for cosine (you can find an explanation here: http://www.intmath.com/analytic-trigonometry/2-sum-difference-angles.php ) \(\cos(v-u)=\cos{v} \cos{u} + \sin{v} \sin{u} \) Plug in the values that you are given, \(\cos{v}=-\dfrac{4}{5}\) and \(\sin{u}=-\dfrac{5}{13}\) Then using special right triangles, you can figure out the two remaining quantities, \(\cos{u}\) and \(\sin{v}\)

OpenStudy (a1234):

Where's the mistake?

OpenStudy (displayerror):

We wouldn't know, you didn't show your work.

jimthompson5910 (jim_thompson5910):

yes please post your work/steps

OpenStudy (a1234):

cosvcosu + sinvsinu -4/5(12/13) + 3/5(5/13) -48/65 + 15/65 = -33/65

jimthompson5910 (jim_thompson5910):

since angle u is in the second quadrant, cos(u) should be negative

OpenStudy (displayerror):

I think sin u can be negative

jimthompson5910 (jim_thompson5910):

sin(u) is positive

jimthompson5910 (jim_thompson5910):

sin(u) = 5/13 cos(u) = -12/13 sin(v) = 3/5 cos(v) = -4/5

OpenStudy (a1234):

-4/5(-12/13) = 48/65 48/65 + 15/65 = 63/65

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (a1234):

Thanks to both of you!

jimthompson5910 (jim_thompson5910):

btw this is extra trivia I guess but cos(v - u) = cos(u-v) because cos(u-v) = cos(-1*(-u+v)) cos(u-v) = cos(-(v-u))) cos(u-v) = cos(v-u) ... using the property that cos(-x) = cos(x) and you're welcome